OpenStudy (anonymous):

The top of a ladder slides down a vertical wall at a rate of 0.125 m/s. At the moment when the bottom of the ladder is 5 m from the wall, it slides away from the wall at a rate of 0.3 m/s. How long is the ladder?

OpenStudy (anonymous):

Ok so I know that I have to set it up as 2x(dx/dt)+2y(dy/dt)=2z(dz/dt)

OpenStudy (anonymous):

But how do I find z with just the information given?

OpenStudy (perl):

correct so far

OpenStudy (perl):

you can use pyth. theorem x^2 + y^2 = z^2

OpenStudy (anonymous):

But I only have length of one side

OpenStudy (perl):

z is constant, so dz/dt = 0

OpenStudy (perl):

z is the length of the ladder, that doesnt change

OpenStudy (perl):

so 2x(dx/dt)+2y(dy/dt)=2z(dz/dt) becomes 2x(dx/dt)+2y(dy/dt)=2z*0 2x(dx/dt)+2y(dy/dt)=0

OpenStudy (anonymous):

oh that makes a lot of sense thanks!

OpenStudy (perl):

and divide out 2